SURFACE ELECTRON EJECTION BY LASER EXCITED METASTABLES SPECTROSCOPY OF ACETY-LENE

\underline{R. L. THOM}, A. P. MISHRA, and R. W. FIELD, Department of Chemistry and G. R. Harrison Spectroscopy Laboratory, Massachusetts Institute of Technology, Cambridge, MA 02139.

We report recent progress in attempts to observe rotationally resolved spectra of directly laser-excited vibrational levels of the third triplet excited electronic state (T_{3}) of acetylene, detected via Auger electron ejection from a gold metal surface. Preliminary results obtained with narrow linewidth ( 0.005 cm^{−1}) excitation using a frequency-doubled pulse-amplified cw ring laser system will be displayed. Insights into singlet-triplet interaction in acetylene gained from such high-resolution simultaneous laser-induced fluorescence and surface electron ejection by laser excited metastables spectroscopy will be discussed.